OpenStudy (anonymous):

restrict the domain of the sine function so that the inverse is a function. Which description best describes how she could restrict the domain?

OpenStudy (anonymous):

A) So that y = sin(x) is always increasing. B) So that y = sin(x) only has one maximum. C) So that y = sin(x) only has one minimum. D) So that y = sin(x) only has two maximums.

OpenStudy (anonymous):

@Hero @pooja195 @Ashleyisakitty @Kainui

OpenStudy (kmeis002):

Do you know the property a function \(f\) must have to be invertible?

OpenStudy (anonymous):

No, i dont remember it

OpenStudy (kmeis002):

Ah, alright. Do you know what property makes a function a function, that is when is a relationship/graph not a function?

OpenStudy (kmeis002):

If we know this, we can apply it to the inverse and come up with a rule.

OpenStudy (anonymous):

I thought it you got a value from plugging it into it then it was a function

OpenStudy (anonymous):

thought if*

OpenStudy (kmeis002):

Hmm, possible depending on what you mean. If you plug in a value and you know the output with certainty, than it is a function. That is, for every input there is a unique output. Often times, teachers call this passing the "vertical line" test.